# Parametric-Task MAP-Elites

### v1

@@ Abstract.
Optimizing a set of functions simultaneously by leveraging their similarity is called multi-task optimization. Current black-box multi-task algorithms only solve a finite set of tasks, even when the tasks originate from a continuous space. In this paper, we introduce Parametric-task MAP-Elites (PT-ME), a novel black-box algorithm to solve continuous multi-task optimization problems. This algorithm (1) solves a new task at each iteration, effectively covering the continuous space, and (2) exploits a new variation operator based on local linear regression. The resulting dataset of solutions makes it possible to create a function that maps any task parameter to its optimal solution. We show on two parametric-task toy problems and a more realistic and challenging robotic problem in simulation that PT-ME outperforms all baselines, including the deep reinforcement learning algorithm PPO.
@@
copyright: noneconference: PREPRINT; February; 2024doi:
@@ 1. Introduction
Optimization is ubiquitous in today’s world (Fletcher, 1987). While many problems require optimizing only one cost function, many involve a set of related functions, each representing a different “task”. For example, in robotics, it can be useful to learn walking gaits for various legged-robot morphologies (Mouret and Maguire, 2020), with each task corresponding to maximizing the walking speed for one morphologie. Instead of solving each optimization individually, simultaneously solving all the tasks should decrease the total computation cost. Another illustration can be found in machine learning, where optimizing the hyperparameters of an algorithm for specific datasets can make a difference; here, each task is the minimization of the loss function for each dataset (Huang et al., 2021).
@@
To our knowledge, all black-box multi-task algorithms solve a finite set of tasks, whereas many problem specifications are continuous. This means that continuous task spaces have to be discretized. For example, in robotics, the morphology is characterized by continuous length of limbs and joint bounds, and any discretization is arbitrary. When the cost function is differentiable, a few algorithms exist (i.e., Parametric programming (Barnett, 1968; Gal and Nedoma, 1972; Pistikopoulos et al., 2000; Fiacco, 1976; Pappas et al., 2021)) to solve the continuous multi-task problem, which we call the parametric-task optimization problem, but we are not aware of any black-box algorithm for this case.
The contribution of this paper is a black-box parametric-task algorithm called Parametric-Task MAP-Elites (PT-ME) inspired by Multi-Task MAP-Elites (MT-ME) (Mouret and Maguire, 2020), an elitist multi-task algorithm. Our central idea is to design an algorithm that solves as many tasks as possible within its allocated budget, hence asymptotically solving the parametric-task optimization problem by filling the task space with high-quality solutions. The resulting dense dataset can then be distilled, for instance, with a neural network, making it possible to generalize and obtain a high-quality solution for any task.
@@
We evaluate our method and its ablations extensively on two parametric-task optimization toy problems, 10-DoF Arm (used in (Mouret and Maguire, 2020; Wang et al., 2022; Huynh Thi Thanh et al., 2023a)) and Archery (a more challenging problem with sparse rewards), and a more realistic parametric-task optimization, Door-Pulling, consisting of a humanoid robot pulling a door open. We compare against several baselines and show that our method effectively solves the parametric-task optimization problem while being more efficient than all baselines and ablations.
@@ 3.1. Parametric Programming
Optimizing a function parameterized with discrete or continuous parameters is studied in mathematics and is called parametric programming (or multiparametric programming for multiple parameters) (Barnett, 1968; Gal and Nedoma, 1972; Pistikopoulos et al., 2000; Fiacco, 1976; Pappas et al., 2021). The main algorithms divide the parameter space into critical regions and solve the optimization problem on each critical region using classical mathematical tools, i.e., Lagrange multipliers and Karush-Kuhn-Tucker conditions. There are three families of methods: multiparametric linear programming (mpLP) (Gal and Nedoma, 1972), multiparametric quadratic programming problems (mpQP) (Pistikopoulos et al., 2000), and multiparametric nonlinear programming (Fiacco, 1976), which approximates nonlinear functions.
@@ 3.2. Multi-Task Optimization
A handful of works tackle more than a dozen tasks: 30 in (Liaw and Ting, 2017, 2019), 50 in (Liang et al., 2022), 500 in (Wang et al., 2022), and 2000 in (Huynh Thi Thanh et al., 2023b). In this paper, we take inspiration from MT-ME (Mouret and Maguire, 2020), which scales up to 5000 tasks.
@@ 3.3. Multi-Task MAP-Elites
MT-ME (Mouret and Maguire, 2020) has applications in robotics, e.g., instead of optimizing each morphology individually, it simultaneously searches for optimal walking gaits for $2000$ morphologies. The resulting walking gaits lead to distances the robots reach three times superior to CMA-ES (Hansen et al., 2003) run individually with the same number of overall iterations. It has applications in industrial scenarios, such as optimizing ergonomic behaviors for different worker morphologies in an industrial workstation (Zhong et al., 2023). Multi-Task Multi-Behavior MAP-Elites (Anne and Mouret, 2023) is an extension that solves the multi-task Quality-Diversity problem and outperforms MAP-Elites (Mouret and Clune, 2015) run on each task individually.
@@ 4.3. Variation Operator 2: Linear Regression
    looks for the corresponding cell $c_{i}$ (L.36) (i.e., the closest centroid $\theta_{c}$ using a KDTree (Maneewongvatana and Mount, 1999) with the SciPy implementation (Virtanen et al., 2020) precomputed at the creation of the archive (L.15));
@@
    looks $c_{i}$’s adjacent cells $C_{adj}$ (using a Delaunay triangulation (Delaunay, 1928) with the SciPy implementation (Virtanen et al., 2020) precomputed at the creation of the archive (L.14)) and extract their task parameters $\boldsymbol{\theta}$ and solutions $\boldsymbol{x}$ (L.37-38);
@@ 4.4. PT-ME Algorithm
Algorithm 1 details PT-ME’s pseudo-code [^1] Source code: hidden for anonymity during review. The two major changes with MT-ME (Mouret and Maguire, 2020) are highlighted in colors and summarized to (1) a uniform sampling at each iteration of a new task parameter on which to evaluate a candidate solution and (2) a new variation operator using local linear regression used for 50% of the iterations.
@@
19:     $A[\theta_{c}].x=\texttt{random.uniform}(0,1,d_{x})$triangleright $x:$ the solution
20:     $A[\theta_{c}].f=\texttt{fitness}(A.t,A.x)$triangleright $f$: the fitness
@@
44:          $A[c_{i}].\theta=\theta$
45:          $A[c_{i}].x=x$
46:          $A[c_{i}].f=f$
47:         if tournament was used thennormal-$\triangleright$ Update the bandit
@@ 5.1. Considered Domains
This toy problem is presented in MT-ME (Mouret and Maguire, 2020) and reused in (Wang et al., 2022; Huynh Thi Thanh et al., 2023a). It consists in finding the joint positions for different 10-DoF arm morphologies that put the end effector as close to a fixed target (Fig. 2.a).
@@ 5.1.3. Door-Pulling
The WBC solves at 1000Hz a quadratic programming problem that optimizes the robot joint positions to satisfy the quasi-static dynamics while trying to satisfy the top-level commands (Dalin et al., 2021).
@@ 5.2.1. Baselines
  • CMA-ES iterates CMA-ES (Hansen et al., 2003) for random task parameters till it reaches the budget $B$ (“How good is CMA-ES?”);
  • CMA-ES (max 10) does the same but enforces a maximum of 10 iterations by task parameter (“Does CMA-ES gain from seeing more different tasks?”);
@@
  • MT-ME runs MT-ME (Mouret and Maguire, 2020) with $5000$ fixed tasks sampled with CVT (Du et al., 1999) (“Do we need a new algorithm?”).
@@ 5.2.3. Evaluations and Measures
After a budget $B$ of $100000$ evaluations, we want to evaluate both the coverage and the quality of the solutions with the idea that a perfect algorithm should find optimal solutions evenly spread on the task parameter space. We use the QD-Score introduced in (Mouret and Clune, 2015) for a finite set of tasks:
@@
As PT-ME stores all the evaluations from $B$ different tasks (L.42), we recompute the elites for different resolutions (Alg. 2), ranging from only one cell (the best-found solution on all tasks) up to $B$ cells evenly spread on the task parameter space $\Theta$ (we cannot fill more cells than the budget). We attribute the lowest fitness to each empty cell, meaning that looking at the QD-Score for the highest resolution does not give a complete idea of the algorithm’s performance. To simplify the comparison, we introduce the Multi-Resolution QD-Score (MR-QD-Score) as the mean QD-Score for a range of resolutions from 1 to $B$ cells. More formally:
@@
where $A_{1:n}$ are archives with different resolutions. We used $n=50$ resolutions evenly spread in logspace between 1 and $100000$ cells.
@@ 5.3. Solutions Evaluations
CMA-ES allocates too many evaluations for the same task, which prevents it from covering the space and leads to a lower QD-Score than Random when there are more than 20 cells to fill. Bounding the number of iterations to 10 allows CMA-ES (10 max) to fill more cells with a better solution than Random, but it worsens after 1000 cells. MT-ME finds solutions with quality similar to the best methods and outperforms CMA-ES and CMA-ES (10 max), showing that solving all the tasks together is more efficient than solving them separately. However, due to its fixed set of tasks, it cannot fill more cells in higher resolutions, significantly decreasing its QD-Score after 5000 cells. PPO and all variants of PT-ME have a similar QD-Score profile, succeeding in filling most of the cells with reasonable solutions, even in high resolutions. PT-ME and PPO have the highest MR-QD-Score and are not significantly different.
@@ 5.3.2. Archery
MT-ME outperforms CMA-ES and CMA-ES (10 max), showing that solving all the tasks together is more efficient than solving them separately. For resolutions inferior to $10000$ cells, MT-ME is better than PPO, which fails in this problem due to its poor exploration with a sparse reward.
@@ 5.4. Distillation Evaluations
Figure 5. (a) Inference score, i.e., mean fitness over $10000$ ($1000$ for Door-Pulling) evenly spread new tasks of PT-ME distillation, MT-ME distillation, and PPO for the three parametric-task optimization problems (the line is the median of 20 (10 for Door-Pulling) replications, and the shaded area is between the first and third quantiles). (b) Box plots with PT-ME’s best resolution ( *********$=$p-value$<$$0.001$, ******$=$p-value$<$$0.01$).
We compare MT-ME and PT-ME with distillation and PPO inference capacities by evaluating the fitness of their proposed solution for $10000$ ($1000$ for Door-Pulling) new tasks sampled with CVT (Du et al., 1999) (Fig. 5). PT-ME significantly outperforms PPO on all the three problems (p-value$<$$0.001$).
PT-ME inference score first increases as the resolution of the archive increases, showing the need for a large dataset for distillation. The decrease in inference score for large resolutions (i.e., $\geq 10000$ cells) is due to the elites’ decreases in quality as the selection pressure decreases. Running the algorithm for longer would increase the maximal resolutions with high-quality elites, even though for Archery, PT-ME reaches nearly a perfect score (the lowest run had an inference score of $0.997$). For the two toy problems, PT-ME outperforms MT-ME for resolutions superior to 500 cells and reaches nearly perfect scores, showing that it covers the space with better solutions and solves the parametric-task optimization problem. For Door-Pulling, the inference score is significantly lower than the MR-QD-Score due to the chaotic behavior of the contact dynamics of our simulator.
@@ 6. Discussion and Conclusion
Can PT-ME be used for large task parameter spaces? PT-ME discretizes the task space using CVT (Du et al., 1999), which scales up to large dimensions (Mouret, 2023). However, the archive needs more cells to keep a high resolution in large dimensions, which can slow down the convergence. Our new variation operator uses a local linear regression, which could be affected by a larger task space (i.e., the linear model requires the inversion of a matrix). Any other supervised method (e.g., neural networks) can be applied to perform regression.
Can PT-ME be used for large solution spaces? Two solutions exist to leverage gradient information to scale up MAP-Elites algorithms to large solution spaces (e.g., a policy neural network): leveraging differentiable problems by computing a gradient from the fitness function (Fontaine and Nikolaidis, 2021) (which requires differentiable problems) and training a differentiable critic and exploits it (Nilsson and Cully, 2021) (applicable to black-box problems). Such gradient-based variation operators could be incorporated into our parametric-task framework.
@@ Acknowledgements.
This project is supported by the CPER SCIARAT, the CPER CyberEntreprise, the Direction General de l’Armement (convention Inria-DGA “humanoïde résilient”), the Creativ’Lab platform of Inria/LORIA, and the EurROBIN Horizon project (grant number 101070596).

### v2

@@ Abstract.
Optimizing a set of functions simultaneously by leveraging their similarity is called multi-task optimization. Current black-box multi-task algorithms only solve a finite set of tasks, even when the tasks originate from a continuous space. In this paper, we introduce Parametric-Task MAP-Elites (PT-ME), a new black-box algorithm for continuous multi-task optimization problems. This algorithm (1) solves a new task at each iteration, effectively covering the continuous space, and (2) exploits a new variation operator based on local linear regression. The resulting dataset of solutions makes it possible to create a function that maps any task parameter to its optimal solution. We show that PT-ME outperforms all baselines, including the deep reinforcement learning algorithm PPO on two parametric-task toy problems and a robotic problem in simulation.
@@
copyright: acmcopyrightcopyright: noneconference: PREPRINT; April; 2024doi: 10.48550/arXiv.2402.01275
@@ 1. Introduction
Optimization is ubiquitous in today’s world (Fletcher, 1987). While many problems require optimizing one cost function, many involve a set of related functions, each representing a different “task”. For example, in robotics, it can be useful to learn walking gaits for various legged-robot morphologies (Mouret and Maguire, 2020), with each task corresponding to maximizing the walking speed for one morphology. Instead of solving each optimization individually, solving all the tasks simultaneously should decrease the total computation cost. Another illustration can be found in machine learning, where optimizing the hyperparameters of an algorithm for specific datasets can make a difference; here, each task is the loss function minimization for each dataset (Huang et al., 2021).
@@
To our knowledge, all black-box multi-task algorithms solve a finite set of tasks, whereas many problem specifications are continuous. This means that continuous task spaces have to be discretized. For example, in robotics, the morphology is characterized by continuous length of limbs and joint bounds, and any discretization is arbitrary. When the cost function is differentiable, a few algorithms exist (i.e., Parametric programming (Pistikopoulos et al., 2002)) to solve the continuous multi-task problem, which we call the parametric-task optimization problem, but we are not aware of any black-box algorithm to do so.
The contribution of this paper is a black-box parametric-task algorithm called Parametric-Task MAP-Elites (PT-ME) inspired by Multi-Task MAP-Elites (MT-ME) (Mouret and Maguire, 2020), an elitist multi-task algorithm. Our main idea is to design an algorithm that solves as many tasks as possible within its allocated budget, asymptotically solving the parametric-task optimization problem by filling the task space with high-quality solutions. The resulting dense dataset can then be distilled, for instance, with a neural network, making it possible to generalize and obtain a high-quality solution for any task.
@@
We evaluate our method and its ablations extensively on two parametric-task optimization toy problems, 10-DoF Arm (used in (Mouret and Maguire, 2020; Wang et al., 2022; Huynh Thi Thanh et al., 2023)) and Archery (a more challenging problem with sparse rewards), and a more realistic parametric-task optimization, Door-Pulling, consisting of a humanoid robot pulling a door open. We compare against several baselines and show that our method effectively solves the parametric-task optimization problem while being more efficient than all baselines and ablations.
@@ 3.1. Parametric Programming
Optimizing a function parameterized with discrete or continuous parameters is studied in mathematics and is called parametric programming (or multiparametric programming for multiple parameters) (Barnett, 1968; Gal and Nedoma, 1972; Pistikopoulos et al., 2000; Fiacco, 1976; Pistikopoulos et al., 2002). The main algorithms divide the parameter space into critical regions and solve the optimization problem on each critical region using classical mathematical tools, i.e., Lagrange multipliers and Karush-Kuhn-Tucker conditions. There are three families of methods: multiparametric linear programming (mpLP) (Gal and Nedoma, 1972), multiparametric quadratic programming problems (mpQP) (Pistikopoulos et al., 2000), and multiparametric nonlinear programming (Fiacco, 1976), which approximates nonlinear functions.
@@ 3.2. Multi-Task Optimization
A handful of works tackle more than a dozen tasks: 30 in (Liaw and Ting, 2017), 50 in (Liang et al., 2022), 500 in (Wang et al., 2022), and 2 00020002\,0002 000 in (Huynh Thi Thanh et al., 2023). In this paper, we take inspiration from MT-ME (Mouret and Maguire, 2020), which scales up to 5 00050005\,0005 000 tasks.
@@ 3.3. Multi-Task MAP-Elites
MT-ME (Mouret and Maguire, 2020) has applications in robotics, e.g., instead of optimizing each morphology individually, it simultaneously optimizes walking gaits for 2 00020002\,0002 000 morphologies. The resulting walking gaits allow robots to travel three times farther than those evolved with CMA-ES (Hansen et al., 2003) run individually with the same budget of iterations. It has applications in industrial scenarios, such as optimizing ergonomic behaviors for different worker morphologies in an industrial workstation (Zhong et al., 2023). Multi-Task Multi-Behavior MAP-Elites (Anne and Mouret, 2023) is an extension that solves the multi-task Quality-Diversity problem and outperforms MAP-Elites (Mouret and Clune, 2015) run on each task individually.
@@ 4.3. Variation Operator 2: Linear Regression
    looks for the closest centroid $\theta_{c}$ (L.36) using a KDTree (Maneewongvatana and Mount, 1999) with the SciPy implementation (Virtanen et al., 2020) precomputed at the creation of the archive (L.15);
@@
    looks for the adjacent cells $A[\theta_{c}].{adj}$ (using a Delaunay triangulation (Delaunay, 1928) with the SciPy implementation (Virtanen et al., 2020) precomputed at the creation of the archive (L.14)) and extract their task parameters $\boldsymbol{\theta}$ and solutions $\boldsymbol{x}$ (L.37-38);
@@ 4.4. PT-ME Algorithm
Algorithm 1 details PT-ME’s pseudo-code [^1] Source code: https://zenodo.org/doi/10.5281/zenodo.10926438. The two major changes with MT-ME (Mouret and Maguire, 2020) are highlighted in colors and summarized to (1) a uniform sampling at each iteration of a new task parameter on which to evaluate a candidate solution and (2) a new variation operator using local linear regression used for 50% of the iterations.
@@
[…]
20:     $A[\theta_{c}].f=\texttt{fitness}(A[\theta_{c}].x,A[\theta_{c}].\theta)$triangleright $f$: the fitness
@@
44:          $A[\theta_{c}].\theta=\theta$
45:          $A[\theta_{c}].x=x$
46:          $A[\theta_{c}].f=f$
47:         if tournament was used $\triangleright$ Update the bandit
@@ 5.1. Considered Domains
This toy problem is presented in MT-ME (Mouret and Maguire, 2020) and reused in (Wang et al., 2022; Huynh Thi Thanh et al., 2023). It consists in finding the joint positions for different 10-DoF arm morphologies that put the end effector as close to a fixed target (Fig. 2.a).
@@ 5.1.3. Door-Pulling
The WBC solves at 1 00010001\,0001 000 Hz a quadratic programming problem that optimizes the robot joint positions to satisfy the quasi-static dynamics while trying to satisfy the top-level commands (Dalin et al., 2021).
@@ 5.2.1. Baselines
  • CMA-ES iterates CMA-ES (Hansen et al., 2003) on a random task parameter till it reaches a stopping criterium, then switches to another till it reaches the budget $B$ (“How good is CMA-ES?”);
  • CMA-ES (max 10) does the same but enforces a maximum of 10 iterations per task parameter before switching to another (“Does CMA-ES gain from seeing more different tasks?”);
@@
  • MT-ME runs MT-ME (Mouret and Maguire, 2020) with 5 00050005\,0005 000 fixed tasks sampled with CVT (Du et al., 1999) (“Do we need a new algorithm?”).
@@ 5.2.3. Evaluations and Measures
After a budget $B$ of 100 000100000100\,000100 000 evaluations, we want to evaluate both the coverage and the quality of the solutions with the idea that a perfect algorithm should find optimal solutions evenly spread on the task parameter space. We use the QD-Score introduced in (Mouret and Clune, 2015) for a finite set of tasks:
@@
As PT-ME stores all the evaluations $E$ from $B$ different tasks (L.42), we recompute the elites for different resolutions, ranging from one cell (the best-found solution on all tasks) up to $B$ cells evenly spread on the task parameter space $\Theta$ (we cannot fill more cells than the budget). We attribute the lowest fitness to each empty cell, meaning that looking at the QD-Score for the highest resolution does not give a complete idea of the algorithm’s performance. To simplify the comparison, we introduce the Multi-Resolution QD-Score (MR-QD-Score) as the mean QD-Score for archives recomputed using Alg. 2 with a range of resolutions from 1 to $B$ cells. More formally:
@@
where $N_{1:n}$ are different resolutions. We used $n=50$ resolutions evenly spread in logspace between 1 and 100 000100000100\,000100 000 cells.
@@ 5.3. Solutions Evaluations
CMA-ES allocates too many evaluations for the same task, which prevents it from covering the space and leads to a lower QD-Score than Random when there are more than 20 cells to fill. Bounding the number of iterations to 10 allows CMA-ES (10 max) to fill more cells with a better solution than Random, but it worsens after 1 00010001\,0001 000 cells. MT-ME finds solutions with quality similar to the best methods and outperforms CMA-ES and CMA-ES (10 max), showing that solving all the tasks together is more efficient than solving them separately. However, due to its fixed set of tasks, it cannot fill more cells in higher resolutions, significantly decreasing its QD-Score after 5 00050005\,0005 000 cells. PPO and all variants of PT-ME have a similar QD-Score profile, succeeding in filling most of the cells with reasonable solutions, even in high resolutions. PT-ME and PPO have the highest MR-QD-Score and are not significantly different.
@@ 5.3.2. Archery
MT-ME outperforms CMA-ES and CMA-ES (10 max), showing that solving all the tasks together is more efficient than solving them separately. For resolutions inferior to 10 0001000010\,00010 000 cells, MT-ME is better than PPO, which fails in this problem due to its poor exploration with a sparse reward.
@@ 5.4. Distillation Evaluations
Figure 5. (a) Inference score, i.e., mean fitness over 10 0001000010\,00010 000 (1 00010001\,0001 000 for Door-Pulling) evenly spread new tasks of PT-ME distillation, MT-ME distillation, and PPO for the three parametric-task optimization problems (the line is the median of 20 (10 for Door-Pulling) replications, and the shaded area is between the first and third quantiles). (b) Box plots with PT-ME’s best resolution ( ***$=$p-value$<$$0.001$, **$=$p-value$<$$0.01$).
We compare MT-ME (for one resolution) and PT-ME (for different resolutions) with distillation and PPO inference capacities by evaluating the fitness of their proposed solution for 10 0001000010\,00010 000 (1 00010001\,0001 000 for Door-Pulling) new tasks sampled with CVT (Du et al., 1999) (Fig. 5). PT-ME significantly outperforms PPO on all the three problems (p-value$<$$0.001$).
PT-ME inference score first increases as the resolution of the archive increases, showing the need for a large dataset for distillation. The decrease in inference score for large resolutions (i.e., 10 000absent10000$\geq$ 10\,000≥ 10 000 cells) is due to the elites’ decreases in quality as the selection pressure decreases. Running the algorithm for longer would increase the maximal resolutions with high-quality elites, even though for Archery, PT-ME reaches nearly a perfect score (the lowest run had an inference score of $0.997$). For the two toy problems, PT-ME outperforms MT-ME for resolutions superior to 500 cells and reaches nearly perfect scores, showing that it covers the space with better solutions and solves the parametric-task optimization problem. For Door-Pulling, the inference score is significantly lower than the MR-QD-Score due to the chaotic behavior of the contact dynamics of our simulator.
@@ 6. Discussion and Conclusion
Can PT-ME be used for large task parameter spaces? PT-ME discretizes the task space using CVT (Du et al., 1999), which scales up to large dimensions (Mouret, 2023). However, the archive needs more cells to keep a high resolution in large dimensions, which can slow down the convergence. Our new variation operator uses a local linear regression, which could be affected by a larger task space (i.e., the linear model requires the inversion of a matrix). Future work could study other supervised methods (e.g., neural networks) to perform regression.
Can PT-ME be used for large solution spaces? Future work could study the integration of two variation operators to leverage gradient information for scaling up MAP-Elites algorithms to large solution spaces (e.g., a policy neural network): leveraging differentiable problems by computing a gradient from the fitness function (Fontaine and Nikolaidis, 2021) (which requires differentiable problems) and training a differentiable critic and exploits it (Nilsson and Cully, 2021) (applicable to black-box problems). Such gradient-based variation operators could be incorporated into our parametric-task framework.
@@ Acknowledgements.
This project is supported by the CPER SCIARAT, the CPER CyberEntreprises, the Direction General de l’Armement (convention Inria-DGA “humanoïde résilient”), the Creativ’Lab platform of Inria/LORIA, the EurROBIN Horizon project (grant number 101070596), the Agence de l’Innovation de Défense (AID), and the ANR in the France 2030 program through project PEPR O2R AS3 (ANR-22-EXOD-007).